It is reported in USA Today that the average flight cost nationwide is $484.314. You have never paid close to that amount and you want to perform a hypothesis test that the true average is actually greater than $484.314. The hypotheses for this situation are as follows: Null Hypothesis: μ ≤ 484.314, Alternative Hypothesis: μ > 484.314. A random sample of 34 flights shows an average cost of $489.228 with a standard deviation of $53.5284. What is the test statistic and p-value for this test?

1) Test Statistic: 0.535, P-Value: 0.298

2) Test Statistic: -0.535, P-Value: 0.702

3) Test Statistic: -0.535, P-Value: 0.298

4) Test Statistic: 0.535, P-Value: 0.596

5) Test Statistic: 0.535, P-Value: 0.702

It is reported in USA Today that the average flight cost nationwide is $353.57. You have never paid close to that amount and you want to perform a hypothesis test that the true average is actually less than $353.57. The hypotheses for this situation are as follows: Null Hypothesis: μ ≥ 353.57, Alternative Hypothesis: μ < 353.57. You take a random sample of national flight cost information and perform a one sample mean hypothesis test. You observe a p-value of 0.8469. What is the appropriate conclusion? Conclude at the 5% level of significance.

1) The true average flight cost is significantly less than $353.57.

2) We did not find enough evidence to say the true average flight cost is greater than $353.57.

3) We did not find enough evidence to say a significant difference exists between the true average flight cost and $353.57.

4) The true average flight cost is greater than or equal to $353.57.

5) We did not find enough evidence to say the true average flight cost is less than $353.57.
Answer #1

Q 1) The test statistic is

t = \frac{\overline{x} -\mu }{s/\sqrt{n}} = \frac{489.228-484.314}{53.5284/\sqrt{34}} =0.535

The P-Value is .298

Answer:  1) Test Statistic: 0.535, P-Value: 0.298

Q 2) Since p value > significance level. we fail to reject H0.

Answer: 5) We did not find enough evidence to say the true average flight cost is less than $353.57.